OpenStudy (darkbluechocobo):

Use elimination to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) 2x + y = 12 -5x - y = -33

OpenStudy (***[isuru]***):

add them together

OpenStudy (***[isuru]***):

wt will u get ?

OpenStudy (darkbluechocobo):

-3x=-21

OpenStudy (***[isuru]***):

go on..

OpenStudy (darkbluechocobo):

x=7

OpenStudy (***[isuru]***):

correct

OpenStudy (***[isuru]***):

so...wt is "y" ?

OpenStudy (darkbluechocobo):

-5x+33=y

OpenStudy (darkbluechocobo):

would that work?

OpenStudy (darkbluechocobo):

-35+33=y -2=y

OpenStudy (***[isuru]***):

that's correct bro..

OpenStudy (darkbluechocobo):

(7,-2) :D and the last one Use elimination to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) 5x - 3y = 42 8x + 5y = 28

OpenStudy (darkbluechocobo):

13x+2y=70

OpenStudy (***[isuru]***):

nope.. don't add them together

OpenStudy (darkbluechocobo):

Oh substitution?

OpenStudy (***[isuru]***):

no substitution either... questions asks for elimination

OpenStudy (darkbluechocobo):

*facepalm* >.<

OpenStudy (***[isuru]***):

multiply first one by 5

OpenStudy (***[isuru]***):

25 x - 15y = 210 ----(1)

OpenStudy (***[isuru]***):

multiply the second by 3

OpenStudy (***[isuru]***):

24x + 15y = 84 ---(2)

OpenStudy (***[isuru]***):

now (1) + (2)

OpenStudy (darkbluechocobo):

3 o.o?

OpenStudy (***[isuru]***):

5x - 3y = 42 ----(1) (1) x 5 25 x - 15y = 210 ----(2) 8x + 5y = 28 ----(3) (3) x 3 24x + 15y = 84 ---(4) (2) + (4)

OpenStudy (darkbluechocobo):

6

OpenStudy (***[isuru]***):

yep

OpenStudy (darkbluechocobo):

(3,6)?

OpenStudy (***[isuru]***):

x = 6 y = -4 (6 , -4)

OpenStudy (darkbluechocobo):

o.o Can you explain what you just did to me cause elimination I do not understand

OpenStudy (***[isuru]***):

in elimination ... u have to eliminate a term in order to find out the other one

OpenStudy (***[isuru]***):

our 2 equations were 5x - 3y = 42 8x + 5y = 28 we could use addition or subtraction and sometimes division for the elimination..

OpenStudy (***[isuru]***):

but here we can't do any of them cause the values of "x" an "y" r different in 2 equations,,, but if we somehow manage to transform the 2 equations in a way such that there will be one term in the same value then we could use add or sub

OpenStudy (***[isuru]***):

so ... i followed these steps to get 15y in both equations 5x - 3y = 42 ----(1) (1) x 5 25 x - 15y = 210 ----(2) 8x + 5y = 28 ----(3) (3) x 3 24x + 15y = 84 ---(4) 'cause then i can use addition to eliminate y got it ?

OpenStudy (***[isuru]***):

still confused ??

OpenStudy (darkbluechocobo):

I see

OpenStudy (darkbluechocobo):

It does make sense

OpenStudy (***[isuru]***):

so... thumbs up huh ?

OpenStudy (darkbluechocobo):

Yes :p Thank you sir

OpenStudy (darkbluechocobo):

You helped out so much :3

OpenStudy (***[isuru]***):

u r welcome bro...

OpenStudy (darkbluechocobo):

Thank you again I can't describe how much I thank you

OpenStudy (***[isuru]***):

haha... lol then u r welcome again too bro...